at 37 degrees Celsius, 0.30M sucrose has about the same osmotic pressure as blood. What is the osmotic pressure of blood?
1
Expert's answer
2020-09-21T06:40:14-0400

We know the osmotic pressure of a nonvolatile nonelectrolyte solution is

"\Pi" =CRT

C=molar concentration of electrolyte=0.30(M),

R=universal gas constant=0.082L-atm/mol k

T=kelvin temperature=37+273=310K

Putting all these value in above equation we get"\Pi" =(0.30*0.082*310)atm=7.62atm,as the blood has same osmotic pressure with sucrose solution the osmotic pressure of blood is 7.62atm